The City Cab Co. charges $2.10 for the first 1/5 of a mile and $0.15 for each additional 1/10 of a mile. The cost of a 3 mile ride is:
Answer by rfer(16322)

30/10-2/10=28/10
28/10*0.15=4.20
4.20+2.10=$6.30